# Evidence of mini-jet emission in a large emission zone from a magnetically-dominated gamma-ray burst jet

### v2

@@ Abstract
The second brightest GRB in history, GRB230307A provides an ideal laboratory to study the details of GRB prompt emission thanks to its extraordinarily high photon statistics and its single broad pulse overall shape characterized by an energy-dependent fast-rise-exponential-decay (FRED) profile. Here we demonstrate that its broad pulse is composed of many rapidly variable short pulses, rather than being the superposition of many short pulses on top of a slow component. Such a feature is consistent with the picture of many mini-jets due to local magnetic reconnection events in a large emission zone far from the GRB central engine, as envisaged in the internal-collision-induced magnetic reconnection and turbulence (ICMART) model, but raises a great challenge to the internal shock models that attribute all variability components to collisions among different shells. Since relativistic mini-jets demand strong magnetization in the outflow, this work provides strong evidence for a Poynting-flux-dominated jet composition of this bright GRB.
Key Laboratory of Particle Astrophysics, Institute of High Energy Physics, Chinese Academy of Sciences, Beijing 100049, China.
University of Chinese Academy of Sciences, Chinese Academy of Sciences, Beijing 100049, China.
Department of Astronomy, Beijing Normal University, Beijing 100875, People’s Republic of China.
Nevada Center for Astrophysics, University of Nevada Las Vegas, NV 89154, USA.
Department of Physics and Astronomy, University of Nevada Las Vegas, NV 89154, USA.
ICRANet, Piazza della Repubblica 10, I-65122 Pescara, Italy
ICRA, Dipartamento di Fisica, Sapienza Università di Roma, Piazzale Aldo Moro 5, I-00185 Rome, Italy
INAF, Osservatorio Astronomico d’Abruzzo, Via M. Maggini snc, I-64100, Teramo, Italy
College of Electronic and Information Engineering, Tongji University, Shanghai 201804, China
College of Physics and Hebei Key Laboratory of Photophysics Research and Application, Hebei Normal University, Shijiazhuang, Hebei 050024, China
Guizhou Provincial Key Laboratory of Radio Astronomy and Data Processing, Guizhou Normal University, Guiyang 550001, China
School of Physics and Optoelectronics, Xiangtan University, Yuhu District, Xiangtan, Hunan, 411105, China
School of Computer and Information, Dezhou University, Dezhou 253023, China
The origin of the prompt emission of gamma-ray bursts (GRBs) is still subject to debate because of the not-well-constrained jet composition, location of the emission region, and mechanism with which -rays are produced[1]. Depending on the unknown jet composition, three emission sites are commonly discussed in GRB prompt emission models. For a matter-dominated fireball, the observed emission is likely a superposition of a thermal component originating from the fireball photosphere at cm[2, 3, 4] and a non-thermal component orignating from synchrotron-radiating electrons accelerated from internal shocks at cm[5, 6, 7]. Within these models, rapid variability in GRB prompt emission lightcurves is attributed to erratic central engine activities, and the emission radius is no larger than the internal shock radius defined by , where is the variability timescale of the short pulses, and is the average bulk Lorentz factor of the jet. The non-detection of bright photosphere emission[8, 9] and non-detection of high-energy neutrinos from any GRBs[10, 11] strongly suggested that a large fraction of GRBs do not have a matter-dominated composition. Rather, the outflows are likely Poynting-flux-dominated even at radii way above the photosphere. Within such a picture, the emission from the photopshere is suppressed. If the magnetization factor is in the emission region, mini-jets may be emitted[12, 13, 14] because magnetic reconnection events would drive a local relativistic flow with Lorentz factor . A well-developed model in this regime is the internal-collision-induced magnetic reconnection and turbulence (ICMART) model[14], which envisages an emission radius cm, where is the timescale of the broad pulses in GRB lightcurves. Within this model[15, 16], short pulses with timescale do not correspond to individual shell collisions, but are rather produced by individual reconnection events occuring at a large emission site defined by . The emission in the timescale of is expected to follow the single-shell evolution and displays an energy-dependent behavior such as energy-dependent duration, spectral lags and global hard-to-soft evolution[17, 18]. The emission in the timescale of , on the other hand, reflects the behavior of individual reconnection events, and is not expected to display all those features.
@@
The extremely bright GRB 230307A was firstly reported by the Gravitational wave high-energy electromagnetic counterpart all-sky monitor (GECAM)[19] with trigger time of 15:44:06.650 UT on 7 March 2023 (denoted as T). The Lobster Eye Imager for Astronomy (LEIA, the pathfinder of the Einstein Probe mission) also observed its prompt emission in the soft X-ray band[20]. While the burst’s duration (measured as ) of about 41 seconds aligns with the category of long-duration GRBs, the association of a kilonova[21] and its unique properties in the prompt emission strongly suggest its origin from a compact binary merger event[20].
Despite of the extraordinary brightness, unsaturated data record by GECAM enable us to accurately characterize the temporal properties of GRB 230307A (see Methods). The GECAM time-binned net count rate as a function of time in different energy bands (multi-band light curves hereafter) show a typical FRED profile (Figure 1a). Therefore, we use a conventional parameterization of the FRED profile [22] (“Norris05" hereafter)
to fit the light curves in each energy bands. In the above formulation, is the starting instance of the pulse, and are the rising and decaying time scales, respectively. The fitting results are listed in Table 1. The peak time and the width of the pulse are therefore defined as: and . It can be seen in Figure 2 that there is a clear energy dependence in both and , as commonly observed in GRB broad pulses. In the energy range from 6 keV to 300 keV, the and relations are both in a power law function, with the power index -0.38 and -0.37, respectively. Above keV, the energy dependence on both and saturate. This highly correlated and relationship implies self-similar profiles across different energy bands, a feature that has also been reported for other bursts in the literature[22, 23, 24, 25]. For the details of the broad feature fitting see Methods.
[…]
a, The black histograms are 1 s binned net count rates (with background subtracted), whose profile shows a single FRED shape; The gray thin histograms are 0.1 s binned net count rates (with background subtracted); The blue solid curves are best-fit FRED model with the Norris05 formulation, while the blue dashed curves are the best-fit with our new FRED formulation. Gray shadowed regions (the precursor and the dip) are ignored in the fitting. All error bars represent 1 uncertainties of the net count rates.
b, The residuals of the net light curves after fitting with the new FRED formulation.
c, The cross-correlation between the multi-band residuals and that in the 6-30 keV channel. The vertical dotted line indicates the zero lag time.
In order to further demonstrate the self-similar feature of the slow varying profile, we smooth the light curves in a non-parametrical way, and manifest their shape identity after energy-dependent re-scaling in time (see Methods). We stack all the re-scaled and smoothed light curves in the Extended Data Fig 1a. It can be obviously shown that, the shape of the profiles in different energy bands can be "stretched" in time domain into an identical shape with an energy dependent scaling factor. The scaling factor as a function of energy is plotted in the Extended Data Fig 1b.
The self-similarity inspired us to propose another formulation of FRED profile, with one less parameter than that of Norris05:
We fit the multi-band light curves again with the new FRED formulation. The results of fitting are listed in Methods. The sole time scale parameter shows a similar energy dependence on energy, , from 6 to 300 keV, and shows a hint of shallower energy dependence above keV (Figure 2). In the new formulation, one can define peak time as , where we denote the in the channel with . The width of the FRED profile also scales as . Therefore, the found dependence can naturally result in the and relations.
Beside the self-similar FRED profile of the broad feature of GRB0307A’s light curve, the extraordinary brightness of the burst allows us to study the fast varying temporal structures. From the residuals of the light curves (Figure 1b), one can see that there are many rapidly-varying short pulses. One noticeable feature is that even the broad pulse has a clear energy dependence, the short time spikes and dips appear to align at the same instances across different energy bands. We further demonstrate this alignment of the fast temporal features by cross-correlations among multi-band light curve residuals. As shown in the Figure 1c, the peaks are perfectly aligned across the full energy band. This immediately excludes the attempt to explain the self-similar stretchable feature of the multi-band light curves as relativistic time-dilation, as a time-dilation effect would stretch both slow and fast varying pulses together.
We show that some of these prominent short duration structures can be fitted with individual short pulses without an underlying slow component(see Methods). The negative residuals, most prominently the dip at around 18 s in all energy bands, essentially excludes the possibility that short time structures are added onto a broad pulse component. It is also difficult to explain this dip as absorption, since the effective optical depth shows an energy dependence which no known absorption mechanism can reproduce (see Methods). Beside, if we were to explain the dip as absorption, its depth and transience require that the absorbing object has a angular size comparable to that of the emission region, and a very high transverse angular velocity. That will either result in a superluminal transverse velocity of the blocking object, or highly coincidental absorption in our local. The most likely explanation is that, the broad pulse of the light curves is composed of many fast pulses, and the dips are gaps between successive short pulses (see Methods).
The aforementioned results that (a) the overall light curve resembles a classic single pulse, which is wider and peaks later in lower energies in a self-similar way; and that (b) there is no underlying elementary single slow pulse and the apparent broad pulse is composed with many fast pulses***We note that even though the feature (a) has been observed and carefully studied before in several other bursts (e.g. the first pulse of GRB130427A[26]), the feature (b) is ambiguously unveiled for the first time only in this burst., serve as important tests on the GRB prompt emission models. The standard internal shock model attributes short-time pulses as a consequence of collisions of many pairs of shells (see Extended Data Fig. 4 for an illustration). The typical emission region is at[5, 6] from the central engine, where is the bulk Lorentz factor of the GRB, and s is the typical variability time scale of the fast component. Within such a scenario, the broad pulse is defined as the history of the central engine activity. Lightcurves in different bands should share this same history and should not display a global energy-dependent behavior (see Methods: Physical implications for detailed simulations that justify this claim). The observed clear energy-dependence of the broad pulse abandons this interpretation. A modified internal shock model (see Extended Data Fig. 5) may invoke another internal shock at a much larger emission radius defined by , where s is the timescale of the broad pulse. However, within such a picture, the observed emission should be the superposition between the fast and slow components, which requires that fast pulses only grow on top of a background flux defined by the slow component. The very deep, achromatic dip at around 18 s, therefore, rules out such a possibility. We therefore conclude that the internal shock models are clearly disfavored. Since all types of photosphere models invoke emission radius smaller or equal to , they are also disfavored by the same reasoning.
The ICMART model, on the other hand, offers a natural interpretation to the data. Within this model[14, 15, 16], the slow component with a characteristic timescale is related to the typical thickness of the magnetic blobs. The collision site between two magnetically-dominated shells is at . The rapid-varying fast component with timescale , on the other hand, is related to the emission timescale of local mini-jets within the global emission zone (see illustration in Extended Data Fig. 6). The broad pulse is simply the superposition of emission of many mini-jets. No underlying slow component is needed. Significant dips are allowed, especially during the decay phase when high-latitude emission starts to play a role. The location of the dip is consistent with such a phase[20]. Because the fast pulses originate from local mini-jet events, the broad-band emission of the fast component is related to the dynamics of the mini-jets and therefore perfectly aligned in time, as is observed.
The general trend of the energy-dependent peak time and duration is naturally expected within the framework of the ICMART model. Unlike the internal shock model that attributes different short pulses as emission from different emitting fluids, the ICMART model suggests that the emission of the entire broad pulse originates from the same fluid as it expands in space. The magnetic field strength in the emission region decays with radius due to expansion[27], leading to rolling down of the characteristic synchrotron emission frequency with time. As result, higher energy emission peaks earlier and lasts shorter than lower energy emission[17, 18]. The general trend of model prediction matches the observations well[17] (see Methods: Physical implications for detailed simulations within the framework of the ICMART model that justify our claim), even though the self-similarity revealed from the data may require special model parameters, which may be related to the details of magnetic turbulence and reconnection in the emission region, and thus provide new insights into the physics of gamma-ray bursts. It is interesting to note that within this picture, the observed prompt emission originates from a single episode of central engine activity. The long-duration of the burst is related to the spreading of the magnetic blob at large radii. This effect reduces the intrinsic duration of the central engine activity, which makes the burst more consistent with the merger origin of GRB 230307A[28, 20, 29].
1 Method
1.1 Observations
@@
GECAM is a dedicated all-sky gamma-ray monitor constellation funded by the Chinese Academy of Sciences, and now consists of three telescopes, i.e. GECAM-A and GECAM-B[30] micro-satellties launched together on December 10, 2020, and GECAM-C (also called High Energy Burst Searcher, HEBS)[31] onboard SATech-01 experimental satellite launched on July 27, 2022. There are two kinds of detectors in each GECAM telescope: Gamma-Ray Detectors (GRDs) and Charged Particle Detectors (CPDs). GRDs are the main detector of GECAM, each of which is composed of a scintillator and an array of SiPMs. There are 25 GRDs onboard each of GECAM-A and GECAM-B, and 12 GRDs onboard GECAM-C. GECAM is designed to have the highest time resolution (0.1 s) among all GRB detectors ever flown[32], making it one of the most suitable detectors for conducting the temporal research about GRBs.
GRB 230307A triggered GECAM-B in real-time at 15:44:06.650 UT on 7 March 2023 (denoted as T) and also detected by GECAM-C and other gamma-ray monitors (e.g. Fermi/GBM, Konus-Wind), while GECAM-A was offline at that time. The real-time alert data was transmitted instantly with the Global Short Message Communication of Beidou satellite navigation system[33, 34], and processed by the automatic pipeline of GECAM[35], based on which the extreme brightness of GRB 230307A was firstly reported by GECAM to the community[19], initiating many multi-wavelength follow-up observations[36, 37]. GRB 230307A is specially interesting because of its apparent single FRED shape in the overall light curve, and its high fluence and variability.
@@
Thanks to the dedicated design of instrument[31, 38], neither GECAM-B nor GECAM-C suffered from data saturation during the whole burst of GRB 230307A despite of its extreme brightness[20]. High quality of GECAM data allow us to accurately measure the temporal and spectral properties of GRB 230307A. GRD04 of GECAM-B and GRD01 of GECAM-C are selected for the analysis of light curves because of their smallest incident angle to the direction of GRB 230307A. These two detectors both operate in two readout channels: high gain (HG) and low gain (LG), which are independent in terms of data processing, transmission, and dead-time. At the time of this burst, GECAM-C GRDs have a lower energy detection threshold of about 6 keV (owing to less radiation damage on SiPM) while GECAM-B GRDs have a relatively higher energy detection threshold of about 30 keV. For GRD04 of GECAM-B, the energy range of HG channel data are used from about 30 keV to 300 keV while the energy range of LG channel data are used from about 300 keV to 1000 keV. For GRD01 of GECAM-C, only HG channel data are used with the energy range from 6 keV to 30 keV. Though the response of GRD01 of GECAM-C for 6-15 keV and GRD04 of GECAM-B for 300-700 keV is affected by the electronics, this does not have any effect on the analysis of light curves. The background of GECAM-B is estimated by fitting the data from -50 s to -5 s and +160 s to +200 s with the first order polynomials. The background of GECAM-C is estimated by fitting the data from -20 s to -1 s and +170 s to +600 s with a combination of the first and second-order exponential polynomials[20]. The observed light curves are not directly used to identify features of broad pulse and fast pulses. Those features are identified through jointly fitting the background and the signals using various functional forms to describe pulses (Norris05, new FRED, etc.).
@@
1.2 Time-binned lightcurve and broad pulse fitting
We divide the data into eight energy bands (1st to 8th channels), namely, 6-30 keV, where the data are from GECAM-C; 30-70 keV, 70-100 keV, 100-150 keV, 150-200 keV, 200-300 keV, 300-500 keV and 500-1000 keV, where the data are from GECAM-B. The counts are binned every one second. The multiband count rates are then fitted to both the FRED pulse profile with Norris05 and our new FRED function (equations 1,2) superimposed on a background function. The fittings are done with the maximum likelihood method, which can be expressed as by assuming a Poisson distribution of the observed count rate, where the subscript i runs over all time bins from -4 s to 100 s (excluding 0 s to 1 s and 16 s to 20 s) and background part (as described in Observations part of Methods). The posteriors of the fitted parameters are found with a Monte Carlo Markov Chain (MCMC) method. The fitted parameters are listed in extended data Table 1.
In the Figure 2, we plot the the FRED parameters, namely and of Norris05 profile and of the new FRED profile, as functions of energy. The vertical error bars are the uncertainties of the fitted parameters, and the horizontal error bars indicate the width of each band. We fit power law relations between the FRED parameters and their corresponding energies. The power law fitting include the points from the 1st to the 6th channels (6-300 keV). The power law indices are -0.38, -0.37 and -0.36 for , and relations respectively. The uncertainties of the power index are found with Monte-Carlo samplings of the FRED parameters and from their distributions. The distributions of the FRED parameters are assumed to be Gaussian with their fitted uncertainties, while the distribution of energy is the corresponding energy deposition spectrum in each band.

### v4

@@ Abstract
The second brightest GRB in history, GRB230307A, provides an ideal laboratory to study the mechanism of GRB prompt emission thanks to its extraordinarily high photon statistics and its single episode activity. Here we demonstrate that the rapidly variable components of its prompt emission compose an overall broad single pulse-like profile. Although these individual rapid components are aligned in time across all energy bands, this overall profile conspires to show a well-defined energy-dependent behavior which is typically seen in single GRB pulses. Such a feature demonstrates that the prompt emission of this burst is from many individual emitting units that are casually linked in a emission site at a large distance from the central engine. Such a scenario is in natural consistency with the internal-collision-induced magnetic reconnection and turbulence framework, which invokes many mini-jets due to local magnetic reconnection that constantly appear and disappear in a global magnetically-dominated jet.
1 introduction
The origin of the prompt emission of gamma-ray bursts (GRBs) is still subject to debate because of the not-well-constrained jet composition, location of the emission region, and mechanism with which -rays are produced (Zhang, 2018). Depending on the unknown jet composition, three emission sites are commonly discussed in GRB prompt emission models. For a matter-dominated outflow, the observed emission is likely a superposition of a thermal component originating from the fireball photosphere at cm (Paczynski, 1986; Goodman, 1986; Mészáros & Rees, 2000) and a non-thermal component originating from synchrotron-radiating electrons accelerated from internal shocks at cm (Rees & Meszaros, 1994; Kobayashi et al., 1997; Daigne & Mochkovitch, 1998a). There are also models with sub-photospheric dissipiation (see Rees & Mészáros 2005a).
Alternatively, a GRB jet can be Poynting-flux-dominated which may dissipate its magnetic energy to power radiation. If dissipation occurs at small radii below the photosphere, the magnetic energy dissipation would mostly modify the spectrum of photospheric emission (Rees & Mészáros, 2005b; Thompson et al., 2007; Giannios, 2006; Drenkhahn & Spruit, 2002). However, if the Poynting-flux energy is retained and gets dissipated only at a large-enough radius well above the photosphere, mini-jets can be generated through magnetic reconnection in a high- ( is defined as the ratio between Poynting flux and matter flux) flow with a typical mini-jet Lorentz factor , and synchrotron radiation of the particles accelerated from these mini-jets will power fast-varying pulses in a GRB light curve (Lyutikov & Blandford, 2003; Narayan & Kumar, 2009; Zhang & Yan, 2011). A well-developed model in this regime is the internal-collision-induced magnetic reconnection and turbulence (ICMART) model (Zhang & Yan, 2011), which envisages an emission radius cm.
@@
The extremely bright GRB 230307A was observed by Fermi (Fermi GBM Team, 2023) and the Gravitational wave high-energy Electromagnetic Counterpart All-sky Monitor (GECAM) (Xiong et al., 2023) with trigger time of 15:44:06.650 UT on 7 March 2023. The Lobster Eye Imager for Astronomy (LEIA, the pathfinder of the Einstein Probe mission) also observed its prompt emission in the soft X-ray band (Sun et al., 2024). While the burst’s duration (measured as ) of about 41 seconds aligns with the category of long-duration GRBs, the association of a kilonova (Levan et al., 2023a) and its unique properties in the prompt emission strongly suggest its origin from a compact binary merger event (Sun et al., 2024). As will be demonstrated in this article, intriguing features were discovered in the multi-band light curves of this burst. This provides us with unique evidence that the ICMART-like mechanism was at work in the prompt emission of GRB 230307A.
The paper is organized as follows: In the section Observation, we will describe the observation details of GRB 230307A; then in the following section Analysis of the multi-band light curves, we present the analysis of the multi-band data of the burst, and exhibit the aforementioned features; in the section Physical implication, we discuss in detail why these observational features can shed light on the prompt emission mechanism.
2 Observation
@@
GRB 230307A triggered GECAM-B in real-time at 15:44:06.650 UT on 7 March 2023 (denoted as T0) and also detected by GECAM-C and other gamma-ray monitors (e.g. Fermi/GBM (Fermi GBM Team, 2023), STIX (Xiao & Krucker, 2023)), while GECAM-A was offline at that time111GECAM is a dedicated all-sky gamma-ray monitor constellation funded by the Chinese Academy of Sciences, and now consists of three telescopes, i.e. GECAM-A and GECAM-B (Li et al., 2021) micro-satellties launched together on December 10, 2020, and GECAM-C (also called High Energy Burst Searcher, HEBS) (Zhang et al., 2023) onboard SATech-01 experimental satellite launched on July 27, 2022. There are two kinds of detectors in each GECAM telescope: Gamma-Ray Detectors (GRDs) and Charged Particle Detectors (CPDs). GRDs are the main detector of GECAM, each of which is composed of a scintillator and an array of SiPMs. There are 25 GRDs onboard each of GECAM-A and GECAM-B, and 12 GRDs onboard GECAM-C. GECAM is designed to have the highest time resolution (0.1 s) among all GRB detectors ever flown (Xiao et al., 2022), making it one of the most suitable detectors for conducting the temporal research about GRBs.. The real-time alert data was transmitted instantly with the Global Short Message Communication of Beidou satellite navigation system (Zhao et al., 2021; Guo et al., 2023), and processed by the automatic pipeline of GECAMHuang et al. (2023), based on which the extreme brightness of GRB 230307A was reported by GECAM to the community (Xiong et al., 2023), initiating many multi-wavelength follow-up observations (Bom et al., 2023; Levan et al., 2023b). GRB 230307A is specially interesting because of its apparent single FRED shape in the overall light curve, and its high fluence and variability.
@@
Thanks to the dedicated design of instrument (Zhang et al., 2023; Liu et al., 2021), neither GECAM-B nor GECAM-C suffered from data saturation during the whole burst of GRB 230307A despite of its extreme brightness (Sun et al., 2024). High quality of GECAM data allow us to accurately measure the temporal and spectral properties of GRB 230307A. GRD04 of GECAM-B and GRD01 of GECAM-C are selected for the analysis of light curves because of their smallest incident angle to the direction of GRB 230307A. These two detectors both operate in two readout channels: high gain (HG) and low gain (LG), which are independent in terms of data processing, transmission, and dead-time. At the time of this burst, GECAM-C GRDs have a lower energy detection threshold of about 6 keV (owing to less radiation damage on SiPM) while GECAM-B GRDs have a relatively higher energy detection threshold of about 30 keV. For GRD04 of GECAM-B, the energy range of HG channel data are used from about 30 keV to 300 keV while the energy range of LG channel data are used from about 300 keV to 1000 keV. For GRD01 of GECAM-C, only HG channel data are used with the energy range from 6 keV to 30 keV. Though the response of GRD01 of GECAM-C for 6-15 keV and GRD04 of GECAM-B for 300-700 keV is affected by the electronics, this does not have any effect on the analysis of light curves. The background of GECAM-B is estimated by fitting the data from -50 s to -5 s and +160 s to +200 s with the first order polynomials. The background of GECAM-C is estimated by fitting the data from -20 s to -1 s and +170 s to +600 s with a combination of the first and second-order exponential polynomials (Sun et al., 2024). The observed light curves are not directly used to identify features of the broad pulse and fast pulses. Those features are identified through jointly fitting the background and the signals using various functional forms to describe pulses (see in the following section).
@@
3 Analysis of the multiband light curves
3.1 The Norris-like single pulse overall profile
The unsaturated data record by GECAM enable us to accurately characterize the temporal properties of GRB 230307A. The high time-resolution light curves show many rapidly varying structures in the prompt emission (see thin grey lines in Figure 1). Intriguingly, by comparing the light curves in different energy bands222namely, 6-30 keV, where the data are from GECAM-C; 30-70 keV, 70-100 keV, 100-150 keV, 150-200 keV, 200-300 keV, 300-500 keV and 500-1000 keV, where the data are from GECAM-B, we found that the light curves also have slow varying trends, which are slower at lower energy bands (this can be better seen with lower time-resolution light curves. See the binned light curves in Figure 1).
In order to quantitatively study the slow trends, we attempt to fit the low time-resolution light curves with a conventional parameterisation of the FRED (Fast Rise Exponential Decay) profile (Norris et al., 2005):
In the above formulation, is the starting instance of the profile, and are the rising and decaying time scales, respectively. In the fitting process, we assume that the net count (with background model subtracted) in each bin follows a Poisson probability distribution, with the expected value equal to the parameterized FRED profile. The fittings are done with the maximum likelihood method, which can be expressed as , where the subscript i runs over all time bins from -4 s to 100 s (excluding 0 s to 1 s and 16 s to 20 s) and the part used to estimate the background model. The posteriors of the fitted parameters are found with a Monte Carlo Markov Chain (MCMC) method.
The fitting results are listed in Table 1. The best-fit Norris FRED profiles are plotted with multi-band light curves in Figure 1. The peak time and the width of the are therefore defined as: and . It can be seen in Figure fig:1.5 that there is a clear energy dependence in both and : In the energy range from 6 keV to 300 keV, the and relations are both in a power law function, with the power indices -0.38 and -0.37, respectively. Above keV, the energy dependence on both and saturate. Such a profile-energy dependence was commonly found in single GRB pulses (Norris et al., 2005; Liang et al., 2006; Hakkila et al., 2008; Peng et al., 2012).
3.2 The self-similarity of the overall profile
We would like to further explore the profile-energy dependence. As we found above, the peak time and the width of the profile vary with energy with the same power law index. As a consequence, the overall profile of the light curve (the broad pulse) in each energy band is a time-stretch copy of that in the other bands (we follow the terminology of Ref. Norris et al. (2005) to call it a “self similar profile”). In order to further demonstrate the self-similar feature of the slow varying profile, we smooth the light curves in a non-parametrical way, and manifest their shape identity after energy-dependent re-scaling in time.
The smoothed light curves are convolution between a Gaussian kernel and the original light curves. The sigma of the Gaussian kernel is 1.5 s. It can be obviously shown that, the shape of the profiles in different energy bands can be ”stretched” in time domain into an identical shape with an energy dependent scaling factor. We use the smoothed light curve in the 500-1000 keV band as the template, and we fit the template to light curves in other bands, which are scaled by multiplying scaling factors to the their time argument . We stack all the re-scaled and smoothed light curves in Figure 3. It is intuitively shown that the self-similarity of the broad profile, and this conclusion is FRED formulation independent. The scaling factor as a function of energy is plotted in the Figure 4, which again shows a power law dependence.